OpenStudy (anonymous):

1/9y=-2 I see that the answer is -18 but I don't really understand the formula to get the answer. 1/9y=-2 I see that the answer is -18 but I don't really understand the formula to get the answer. @Mathematics

OpenStudy (amistre64):

there is no formula, there is just algebra

OpenStudy (amistre64):

how do we get y all alone? how do we undo a 1/9 ?

OpenStudy (amistre64):

this is a concept known as inverses, how to undo what was done prior

OpenStudy (anonymous):

thanks Amistre64... I am trying to teach this to my daughter. I can do it in my head but I am having a hard time explaining how I got the answer.

OpenStudy (amistre64):

yeah, its the inverse concept that they are working on. The inverse to left is right, top is bottom, addition is subtraction, division is multiplication. multiply both sides by 9 to undo the division by 9

OpenStudy (amistre64):

\[\frac{1}{9}y=-2\] \[(9)\frac{1}{9}y=-2(9)\] \[y=-18\]

OpenStudy (anonymous):

very well stated. thank you very much.

OpenStudy (amistre64):

your welcome, and good luck with it all ;)

OpenStudy (preetha):

My son had a problem with this concept in algebra as wll. So it really helped to see the same operation performed on both sides, then written out and then simplified. good job amistre.